### v3

# On Discounted Infinite-Time Mean Field Games

Abstract
In this paper, we study the infinite-time mean field games with discounting, establishing an equilibrium where individual optimal strategies collectively regenerate the mean-field distribution. To solve this problem, we partition all agents into a representative player and the social equilibrium. When the optimal strategy of the representative player shares the same feedback form with the strategy of the social equilibrium, we say the system achieves a Nash equilibrium.
We construct a Nash equilibrium using the stochastic maximum principle and infinite-time forward-backward stochastic differential equations(FBSDEs). By employing the elliptic master equations, a class of distribution-dependent elliptic PDEs , we provide a representation for the Nash equilibrium. We prove the Yamada-Watanabe theorem and show the weak uniqueness for infinite-time FBSDEs. And we prove that the solutions to a system of infinite-time FBSDEs can be employed to construct viscosity solutions for a class of distribution-dependent elliptic PDEs.
Keywords. discounted infinite-time mean field games, infinite-time forward-backward equations, weak uniqueness, elliptic master equations.
1 Introduction
The study of mean field games was initiated independently by Lasry-Lions (see [9, 10, 11]) and Huang-Malhamé-Caines [8], which is an analysis of limit models for symmetric weakly interacting $(N+1)$-player differential games. It is noteworthy that current theoretical frameworks are primarily developed for finite-time problems, while infinite-time scenarios remain significantly underdeveloped. We refer the reader to [5] for a comprehensive exposition on this subject.
In this paper, we consider a generalized framework for mean field games, which extends the classical finite-time settings to discounted infinite-time mean field games. Our model differs from the fixed-point problems studied in prior works. In our framework, a representative player interacts with a continuum of other players (also referred to as the population or social equilibrium). The dynamic of the private state of the representative player is given by
$X^{x}_{t}=x+\int_{0}^{t}b(X^{x}_{t},\mu_{s},\beta_{s}^{x})\mathrm{d}s+B_{t}.$ (1.1)
Here $\mu_{t}$ is the population distribution , while $\beta^{x}$ is the optimal strategy of the representative player, derived as the solution to the stochastic optimization problem:
$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right],$ (1.2)
under the constraint
$\begin{cases}\mathrm{d}X_{t}^{x,\beta}=b(X_{t}^{x,\beta},\mu_{t},{\beta}_{t})\mathrm{d}t+\mathrm{d}B_{t},\\ X_{0}^{x,\beta}=x.\end{cases}$ (1.3)
Since the population consists of a multitude of homogeneous individuals, the macroscopic state should satisfy
$X^{\xi}_{t}=X^{x,\beta^{x}}|_{x=\xi},$ (1.4)
where $\mathcal{L}_{\xi}=\mu_{0}$ denotes the initial population distribution. We say the population reaches equilibrium if $\mu_{t}=\mathcal{L}_{X^{\xi}_{t}}.$
This model decouples the micro-level agent from the macro-level societal distribution, enabling interconnected analysis of their evolution. The equilibrium is characterized by two consistency conditions:
  • Individual Rationality: The representative player’s optimal strategy is consistent with the perceived social equilibrium.
  • Macro Consistency: The aggregate distribution generated by all players adopting this strategy must equal the posited social equilibrium.
We will employ the stochastic maximum principle and infinite-time forward-backward stochastic differential equations (FBSDEs) to solve this game-theoretic problem, and derive a representation of the equilibrium strategy via an elliptic-type master equation.
The investigation of general nonlinear BSDEs was pioneered by Pardoux and Peng [13, 14] in the early 1990s, which is now a typical tool in stochastic optimization problems. We shall solve the mean field game problem using the Pontryagin’s maximum principle and infinite-time forward-backward stochastic differential equations. The foundational work in [15] proved the existence and uniqueness of solutions to the infinite-time FBSDEs, and subsequent work in [16] investigated a more general setting and established connections with quasilinear elliptic PDEs. Recently, [1] extended this framework to the Mckean-Vlasov FBSDEs, which play crucial roles in solving equilibrium solutions for mean field games. In this paper, we partition all agents into a representative player and the social equilibrium, and characterize the equilibrium state through the following system of infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.5)
$\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.6)
Here $\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy$ is the generalized Hamiltonian, and $\hat{\alpha}(x,y)$ is the minimizer of $\mathcal{H}(x,\mu,a,y)$ with respect to $a$ when we assume $\mathcal{H}$ is separable in variables $\mu$ and $a$. The solution $X_{t}^{\xi}$ to Equation (1.5) represents the population’s state process, whose law corresponds to the population distribution $\mu_{t}$. And the solution $X_{t}^{x}$ to Equation (1.6) is the state process of the representative agent after solving the optimization problem. Notably, it exhibits the same feedback structure as the population’s state process. To further elucidate the relationship in Equation (1.4 ), we introduce the elliptic-type master equations.
First introduced by Lions in lectures [12], the parabolic-type master equations appeared in the context of the theory of mean field games. This is a time-dependent equation that bears profound connections with finite-time mean field game theory. Essentially, it describes a strategic interaction between a representative player and the collective environment. When the Nash equilibrium exists, the master equation provides a powerful tool to characterize the equilibrium cost and control pattern of this system. We refer the reader to [5, 3, 7] for a comprehensive exposition on the subject. In this paper, we propose the elliptic-type master equations, which explicitly characterizes the feedback forms of both the representative player and the social equilibrium. This equation takes the following form:
$\begin{split}rU(x,\mu)=&H(x,\mu,\partial_{x}U(x,\mu))+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (1.7)
Here $\partial_{x},\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu},\partial_{\tilde{x}\mu}$ are $W_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law. Under the assumption that the master equation (1.7) admits a solution with sufficient regularity, we derive the following representation for Equation (1.5) and (1.6):
$Y_{t}^{\xi}=\partial_{x}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{\xi}=\partial_{xx}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}).$ (1.8)
$Y_{t}^{x}=\partial_{x}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x}=\partial_{xx}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}).$ (1.9)
If we further assume that $\tilde{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\mu,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$. For any fixed finite time $T$, we have
$X_{t}^{x}|_{x=\xi}=X_{t}^{\xi},\quad t\in[0,T].$ (1.10)
This is precisely the relationship expressed in Equation (1.4).
This paper is organized as follows: in section 2, we present the preliminaries of problems in this paper; in section 3, we introduce the infinite-time mean field games with discounting and the definition of Nash equilibrium; in section 4 we characterize the equilibrium states through a system of infinite-time FBSDEs and in section 5 we introduce the elliptic master equation to provide a representation for the Nash equilibrium; in section 6, we prove the Yamada-Watanabe theorem for infinite-time FBSDEs and provide a viscosity solution for distribution-dependent elliptic PDEs by virtue of the class of FBSDEs introduced in section 4.
2 Preliminaries
We will use the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$. Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$, such that it can support any measure on $\mathbb{R}$ with finite second moment.
Let $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ be a copy of the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with corresponding Brownian motion $B^{\prime}$, define the larger filtered probability space by
$\tilde{\Omega}\triangleq\Omega\times\Omega^{\prime},\quad\tilde{\mathcal{F}}\triangleq\mathcal{F}\otimes\mathcal{F}^{\prime}\quad\tilde{\mathbb{F}}=\{\tilde{\cal F}_{t}\}_{t\geq 0}\triangleq\{{\cal F}_{t}\otimes{\cal F}^{\prime}_{t}\}_{t\geq 0},\quad\tilde{\mathbb{P}}\triangleq\mathbb{P}\otimes\mathbb{P}^{\prime},\quad\tilde{\mathbb{E}}\triangleq\mathbb{E}^{\tilde{\mathbb{P}}}$. (2.1)
[…]
Let ${\cal P}\triangleq{\cal P}(\mathbb{R})$ be the set of all probability measures on $\mathbb{R}$ and let ${\cal P}_{p}(p\geq 1)$ denote the set of $\mu\in{\cal P}$ with finite $p$-th moment. For any sub-$\sigma$-field $\cal G$$\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, $\cal G$}-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ . For any $\mu,\nu\in{\cal P}_{p}$, we define the $W_{p}$–Wasserstein distance between them as follows:
$W_{p}(\mu,\nu):=\inf\Big{\{}\big{(}\mathbb{E}[|\xi-\eta|^{q}]\big{)}^{1/q}:~\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big{\}}.$
Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}$-valued adapted stochastic process ($v_{t}$) start from $t_{0}$ such that
$\mathbb{E}\left[\int_{t_{0}}^{\infty}e^{-Kt}|v_{t}|^{2}dt\right]<+\infty.$ (2.2)
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$. For each $\mathcal{F}_{0}$-measurable square integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.3)
It has a unique solution $(X_{t},Y_{t},Z_{t})\in L^{2}_{K}(0,+\infty,\mathbb{R}^{3})$ under the following assumptions:
Assumption 2.1
(i) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
$|B(t,x,y,\mu)$ $-B(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.4)
$\leq\ell(|x-x^{\prime}|+|y-y^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime})).\quad\mathrm{a.s.}$
(ii) There exist constants $0<K<2\kappa$ such that for any $t\geq 0$ and any square integrable random variables $X,X^{\prime},Y,Y^{\prime}$
$\mathbb{E}\left[-K\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (2.5)
$\leq-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}}).$
For the detailed proof, we refer the reader to ([1], Theorem 2.1).
We introduce the Wasserstein space and differential calculus on Wasserstein space. For a $W_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$, its $W_{2}$-Wasserstein derivatives[5](also called Lions-derivative), takes the form $\partial_{\mu}U:(\mu,\tilde{x})\in{\cal P}_{2}\times\mathbb{R}\to\mathbb{R}$ and satisfies:
$U({\cal L}_{\xi+\eta})-U(\mu)=\mathbb{E}\big{[}\langle\partial_{\mu}U(\mu,\xi),\eta\rangle\big{]}+o(\|\eta\|_{2}),\ \forall\ \xi\in\mathbb{L}^{2}(\mathcal{F};\mu),\eta\in\mathbb{L}^{2}(\mathcal{F}).$ (2.6)
Let ${\cal C}^{0}({\cal P}_{2})$ denote the set of $W_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$. For ${\cal C}^{1}({\cal P}_{2})$, we mean the space of functions $U\in{\cal C}^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.6). Let ${\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$. Let ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U,\partial_{xxx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U,\partial_{xx\mu}U,\partial_{x\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}^{d}\times{\cal P}_{2}\times\mathbb{R}$.
Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula[2, 5]. For $i=1,2$, let $dX^{i}_{t}\triangleq b^{i}_{t}dt+\sigma^{i}_{t}dB_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}:={\cal L}_{X^{2}_{t}}$. Fix $T>0$ and let all conditions be restricted to the interval [0,T]. Then we have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big{[}\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\hbox{\rm tr$\,$}\big{(}\partial_{xx}U[\sigma_{t}^{1}]^{2}\big{)}\Big{]}(t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big{(}\tilde{\mathbb{E}}_{{\cal F}_{t}}\big{[}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})[\tilde{\sigma}_{t}^{2}]^{2}\big{]}\Big{)}\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.7)
Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ .
3 Infinite-time mean field games with discounting
In this section, we introduce the infinite-time mean field games with discounting. Let $r>0$ represent the time discount factor and $A\subset\mathbb{R}$ be a convex control space. Define $\mathcal{A}\triangleq L^{2}_{r}(0,\infty,A)$ to be the space of all adimissible controls, and $b,f:\mathbb{R}\times{\cal P}_{2}\times A\rightarrow\mathbb{R}$ are two measurable functions.
We consider a population consisting of a continuum of players, where each individual player strategically interacts with the aggregate distribution formed by all other players to minimize their own cost. Let $\mu_{t}$ denote the population distribution and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ denote the initial state with $\mathcal{L}_{\xi}=\mu_{0}$. The state of the representative player with initial value $x$ is given by
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mu_{t},\beta_{s})\mathrm{d}s+B_{t},$ (3.1)
where $\beta\in\mathcal{A}$ is the strategy remains to be determined. The representative player seeks to minimize the cost
$J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (3.2)
Assuming $\beta^{x}\in\mathcal{A}$ minimizes the cost function, the state process of the representative player is $X^{x,\beta^{x}}_{t}$. Since the representative agent can characterize the strategies of all players in the population, we assert that the following fundamental relationship must hold:
$\mu_{t}=\mathcal{L}_{X^{x,\beta^{x}}_{t}|_{x=\xi}}.$ (3.3)
To solve this mean field game problem, we work under the following assumptions:
Assumption 3.1
(i) $b(x,\mu,a)$ is Lipschitz in $(x,\mu,a)$, and $f(x,\mu,a)$ is of at most quadratic growth in $(x,\mu,a)$. There exists a positive constant $\ell$ such that for any $\mu,\mu^{\prime}\in{\cal P}_{2},x\in\mathbb{R},a\in A$,
$|b(x,\mu,a)-b(x,\mu^{\prime},a)|\leq\ell\mathcal{W}_{2}(\mu,\mu^{\prime}).$ (3.4)
(ii) There exists a constant $\lambda>\ell-r/2$ such that for any $a\in A,\mu\in{\cal P}_{2},x,x^{\prime}\in\mathbb{R}$, it holds that
$(x-x^{\prime})\left(b(x,\mu,a)-b(x^{\prime},\mu,a)\right)\leq-\lambda(x-x^{\prime})^{2}.$ (3.5)
These assumptions jointly ensure that: (i) the state process remains confined within the $L^{2}_{r}$ space, and (ii) the cost functional maintains integrability.
Then we partition all agents into two components: (i) the representative player, who dynamically optimizes their strategy, and (ii) the social equilibrium (or mean field), characterizing the macroscopic state shared by the population. For the social equilibrium, its state is governed by the following SDE:
$X_{t}^{\xi,\alpha}=\xi+\int_{0}^{t}b(X_{s}^{\xi,\alpha},\mathcal{L}_{X_{s}^{\xi,\alpha}},\alpha_{s})\mathrm{d}s+B_{t},$ (3.6)
where $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$. We note that, by assumption 3.1, this SDE has a unique strong solution in $L^{2}_{r}$, see ([1], Proposition 2.2) for more details.
The state of the representative player is governed by
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mathcal{L}_{X_{s}^{\xi,\alpha}},\beta_{s})\mathrm{d}s+B_{t}.$ (3.7)
Here, we also require their control $\beta\in\mathcal{A}.$
The representative player seeks to minimize the cost
$J\left(x,\xi;\alpha,\beta\right)=\mathbb{E}\bigg{[}\int_{0}^{+\infty}\mathrm{e}^{-rt}f\big{(}X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi,\alpha}},\beta_{t}\big{)}\mathrm{d}t\bigg{]}.$ (3.8)
For any $(x,\xi)\in\mathbb{R}\times\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$, we consider the infimum
$V(x,\xi;\alpha)\triangleq\inf_{\beta\in\mathcal{A}}J(x,\xi;\alpha,\beta).$ (3.9)
Definition 3.2
We say a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
$\alpha^{x}\in\arg\min_{\beta\in\mathcal{A}}J(x,\xi_{0};\alpha^{\xi_{0}},\beta).$ (3.10)
When this Nash equilibrium $\alpha^{*}$ exists, we have
$\xi,\xi^{\prime}\in\mathbb{L}^{2}(\mathcal{F}_{0}),\quad\mathcal{L}_{\xi^{\prime}}=\mathcal{L}_{\xi}\quad\Longrightarrow\quad\mathcal{L}_{X_{t}^{\xi,\alpha^{\xi}}}=\mathcal{L}_{X_{t}^{\xi^{\prime},\alpha^{\xi^{\prime}}}},\quad~\mbox{for a.e. $t\geq 0$.}$ (3.11)
Therefore, we can define
$V(x,\mu)\triangleq J(x,\xi_{0};\alpha^{\xi_{0}},\alpha^{x}),\quad\quad\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0},\mu).$ (3.12)
Note that our definition of Nash equilibrium differs from that in [1] and earlier finite-time mean field games. In [1], the mean field game is formulated as a fixed-point problem. The population is assumed to be homogeneous, meaning all agents are identical and thus represented by a single representative player. For a given measure flow $(\mu_{t})_{t\geq 0}$ , the representative player wants to minimize
$J^{\mu}(\alpha)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(t,X_{t},\mu_{t},\alpha_{t})\mathrm{d}t\right],$ (3.13)
under the constraint
$\begin{cases}\mathrm{d}X_{t}=b\left(t,X_{t},\mu_{t},{\alpha}_{t}\right)\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (3.14)
Then we require that the law of $X_{t}$ coincides with $\mu_{t}$, which means we need to find a fixed point.
In this paper, we separate a representative player from the population, who only needs to consider their own optimization problem starting from state $x$. The representative player’s state evolution depends on both their own state $X^{x,\beta}$ and the overall population distribution $\mathcal{L}_{X^{\xi,\alpha}}$. Here we would like to emphasize that since there are a large number of players, any change of a representative player doesn’t impact the measure flow $\mathcal{L}_{X^{\xi,\alpha}}$. Under the existence assumption of the Nash equilibrium $\alpha^{*}$ specified in Definition 3.2, the stochastic dynamics of both the population and representative player are characterized by the following stochastic differential equations:
$\begin{cases}X_{t}^{\xi,\alpha^{*}}=\xi+\int_{0}^{t}b(X_{s}^{\xi,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}},{\alpha}^{*}(X_{s}^{\xi,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}}))\mathrm{d}s+B_{t},\\ X_{t}^{x,\alpha^{*}}=x+\int_{0}^{t}b(X_{s}^{x,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}},{\alpha}^{*}(X_{s}^{x,\alpha^{*}},\mathcal{L}_{X_{s}^{\xi,\alpha^{*}}}))\mathrm{d}s+B_{t}.\end{cases}$ (3.15)
We further assume that $\tilde{b}(x,\mu)\triangleq b(x,\mu,\alpha^{*}(x,\mu))$ is Lipschitz continuous in $(x,\mu)$. For any fixed finite time $T$, see [4, 2] , we have
$X_{t}^{x,\alpha^{*}}|_{x=\xi}=X_{t}^{\xi,\alpha^{*}},\quad t\in[0,T].$ (3.16)
This implies that every sample point from the initial population follows the same evolutionary dynamics as our hypothesized representative player, which justifies the mathematical validity of using a single representative player to characterize the behavior of all individuals in the population. Moreover, when all agents in the population adopt the same strategy as the representative player, their collective behavior precisely generates the aggregate distribution $\mathcal{L}_{X_{t}^{\xi,\alpha^{*}}}$ derived from our solution. This justifies why we refer to $X_{t}^{\xi,\alpha^{*}}$ as the social equilibrium.
4 Connection with infinite-time McKean-Vlasov FBSDEs
In this section, we employ the maximum principle to solve the optimization problem for the representative player and then use the infinite-time McKean-Vlasov FBSDE to construct the optimal strategy for the representative player such that the controls of the representative player and the social equilibrium share the same feedback form. Our derivation is based on the following key assumptions on $b,f$:
Assumption 4.1
(i) $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$. Where $b_{0},f_{0}$ are measurable functions on $\mathbb{R}\times{\cal P}_{2}$, $b_{1},f_{1}$ are measurable functions on $\mathbb{R}\times A$.
(ii) $b,f$ are differentiable with respect to $(x,a)$ and $\partial_{a}b,\partial_{a}f$ are Lipschitz continuous in $(x,a)$.
(iii) $H_{0}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)$ is convex with respect to $(x,a)$. $\min_{a\in A}H_{0}(x,\mu,a,y)$ has a unique minimizer $\hat{\alpha}(x,y)$ which is Lipschitz continuous in $(x,y)$.
4.1 Pontryagin’s maximum principle
Assuming the state of the social equilibrium $X_{t}^{\xi}$ is given, we consider the optimization problem for the representative player, whose state is given by
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mathcal{L}_{X_{s}^{\xi}},\beta_{s})\mathrm{d}s+B_{t}.$ (4.1)
The cost functional takes the form
$J(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\mathrm{d}t\right],$ (4.2)
and the representative player wants to solve the minimization problem
$\inf_{\beta\in\mathcal{A}}J(\beta).$ (4.3)
Let us formally derive the maximum principle for the infinite-time control problem. Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ . Let
$R_{t}=\lim_{\epsilon\to 0}\frac{X_{t}^{x,\beta+\epsilon\gamma}-X_{t}^{x,\beta}}{\epsilon}$ (4.4)
be the variation process. Then it can be shown that $R$ satisfies
$\begin{cases}\mathrm{d}R_{t}&=\left(\partial_{x}b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot R_{t}+\partial_{a}b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot\gamma_{t}\right)\mathrm{d}t,\\ R_{0}&=0.\end{cases}$ (4.5)
The function $\beta\to J(\beta)$ is Gâteaux differentiable in the direction $\beta$ and its derivative is given by
$\left.\frac{d}{d\epsilon}J(\beta+\epsilon\gamma)\right|_{\epsilon=0}=\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(\partial_{x}f(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot R_{t}+\partial_{a}f(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\cdot\gamma_{t}\right)dt\right].$ (4.6)
Define the generalized Hamiltonian
$\mathcal{H}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)-rxy,$ (4.7)
and introduce the adjoint process which is determined by an infinite-time BSDE
$\mathrm{d}Y_{t}^{x,\beta}=-\left(\partial_{x}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})\right)\mathrm{d}t+Z_{t}^{x,\beta}\mathrm{d}B_{t}.$ (4.8)
Applying Itô’s formula to the process $(e^{-rt}R_{t}Y_{t}^{x,\beta})$, by simple computation we can deduce that
$\left.\frac{\mathrm{d}}{\mathrm{d}\epsilon}J(\beta+\epsilon\gamma)\right|_{\epsilon=0}=\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\partial_{a}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})\cdot\gamma_{t}\mathrm{d}t\right].$ (4.9)
Thus when $\beta$ is an optimal admissible control with the associated stochastic processes $(X_{t}^{x,\beta},Y_{t}^{x,\beta},Z_{t}^{x,\beta})$, it holds that
$\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x,\beta})=\min_{a\in A}\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},a,Y_{t}^{x,\beta}).$ (4.10)
Recalling our convexity assumptions on $b,f$ in Assumption (4.1), we know that the representative player’s optimal control $\beta$ takes a feedback form, that is
$\beta_{t}=\hat{\alpha}(X_{t}^{x,\beta},Y_{t}^{x,\beta}).$ (4.11)
Now we consider the following two McKean–Vlasov FBSDEs:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=b(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\partial_{x}\mathcal{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{\xi},Y_{t}^{\xi}),Y_{t}^{\xi})\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (4.12)
$\begin{cases}\mathrm{d}X_{t}^{x}=b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}))\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x},Y_{t}^{x}),Y_{t}^{x})\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.13)
Here (4.12) and (4.13) denote the state processes of social equilibrium and the representative player, respectively. Observe that their admissible controls both take the identical feedback form $\hat{\alpha}(x,y)$. We shall prove that when social equilibrium employ this feedback control, the representative player’s loss function is minimized if they use the same feedback form—thereby constituting a Nash equilibrium.
Theorem 4.2
Let $(b,f)$ be differentiable in $(x,a)$, $\mathcal{H}$ be convex in $(x,a)$. Suppose that $\hat{\alpha}$ is Lipschitz continuous and that both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$. If we denote $\hat{\alpha}(X_{t}^{x},Y_{t}^{x})$ as $\alpha^{*}_{t}$ which is an admissible control in $\mathcal{A}$, then we have
$J({\alpha}^{*})=\min_{\beta\in\mathcal{A}}J(\beta).$ (4.14)
Proof. For an arbitrary admissible control $\beta\in\mathcal{A}$ and its associated process
$X_{t}^{x,\beta}=x+\int_{0}^{t}b(X_{s}^{x,\beta},\mathcal{L}_{X_{s}^{\xi}},\beta_{s})\mathrm{d}s+B_{t},$ (4.15)
we have
$J(\alpha^{*})-J(\beta)=$ $\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})-\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x})\right)\mathrm{d}t\right]$ (4.16)
$-\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t})-b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\right)\cdot Y_{t}^{x}\mathrm{d}t\right]$
$+r\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}(X_{t}^{x}-X_{t}^{x,\beta})\cdot Y_{t}^{x}\mathrm{d}t\right].$
Since $X_{t}^{x},X_{t}^{x,\beta},Y_{t}^{x}$ are all belong to $L^{2}_{r}$, we can find a sequence of $T_{i}\rightarrow\infty$ such that
$\mathbb{E}\left[e^{-rT_{i}}(X_{T_{i}}^{x}-X_{T_{i}}^{x,\beta})\cdot Y_{T_{i}}^{x}\right]\to 0.$ (4.17)
Applying Itô’s formula to $e^{-rT_{i}}(X_{T_{i}}^{x}-X_{T_{i}}^{x,\beta})\cdot Y_{T_{i}}^{x}$ and letting $T_{i}\to\infty$, we obtain that
$\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}(X_{t}^{x}-X_{t}^{x,\beta})\left(\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})\right)\mathrm{d}t\right]$ (4.18)
$=$ $\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\left(-r(X_{t}^{x}-X_{t}^{x,\beta})+b(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t})-b(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t})\right)\cdot Y_{t}^{x}\mathrm{d}t\right].$
According to the convexity and differentiability of $\mathcal{H}$, we have
$\mathcal{H}(X_{t}^{x,\beta},\mathcal{L}_{X_{t}^{\xi}},\beta_{t},Y_{t}^{x})-\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})$ (4.19)
$\geq$ $(X_{t}^{x,\beta}-X_{t}^{x})\cdot\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})+(\beta_{t}-{\alpha}^{*}_{t})\cdot\partial_{a}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x})$
$=$ $(X_{t}^{x,\beta}-X_{t}^{x})\cdot\partial_{x}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},\alpha^{*}_{t},Y_{t}^{x}).$
The last equality follows from the fact that $\alpha^{*}_{t}=\hat{a}(X_{t}^{x},Y_{t}^{x})$ is the minimizer of $\min_{a\in A}\mathcal{H}(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}},a,Y_{t}^{x})$. Combining equations (4.16), (4.18), and (4.19), we obtain
$J(\alpha^{*})-J(\beta)\leq 0$ (4.20)
for all admissible control $\beta$. Thus we complete the proof.
4.2 Solvability of mean field game FBSDEs
In this subsection, we will find sufficient conditions for the existence and uniqueness of solutions to (4.12) and (4.13). Considering the linear case, we assume $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
We require a technical lemma about the Lipschitz and convex property of the minimizer $\hat{\alpha}$, the detailed proof of which can be found in ([1], Lemma 3.1).
Lemma 4.3
Suppose $f_{1}$ is once continuously differentiable in $(x,a)$, $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$. And $f_{1}$ is $\eta$-convex in $a$, which means
$f_{1}(x,a^{\prime})-f_{1}(x,a)-(a^{\prime}-a)\cdot\partial_{a}f_{1}(x,a)\geq\eta|a^{\prime}-a|^{2},\quad~\mbox{for all $x\in\mathbb{R}$.}$ (4.21)
Then it holds that
$|\hat{\alpha}(x,y)-\hat{\alpha}(x^{\prime},y^{\prime})|\leq\frac{\ell}{2\eta}|x^{\prime}-x|+\frac{|b_{3}|}{2\eta}|y^{\prime}-y|,$ (4.22)
and for some $a_{0}\in A$
$|\hat{\alpha}(x,y)|\leq\eta^{-1}(|\partial_{a}f_{1}(x,a_{0})|+|b_{2}y|)+|a_{0}|.$ (4.23)
Furthermore, if $A=\mathbb{R}$ and $\partial_{a}f$ is $\zeta$-Lipchitz in $a$, it follows that
$b_{3}(y^{\prime}-y)\cdot\left(\hat{\alpha}(x,y^{\prime})-\hat{\alpha}(x,y)\right)\leq-\frac{2b_{3}^{2}\eta}{\zeta^{2}}(y^{\prime}-y)^{2}.$ (4.24)
Theorem 4.4
Let $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$. Under the following conditions, Assumption 3.1 and 4.1 are satisfied, and both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
(i) There exists a positive constant $k$ such that $|b_{2}|\leq k$ and $b_{1}\geq k-\frac{r}{2}$. $f_{0}(x,\mu)$ is once continuously differentiable in $x$ and of at most quadratic growth in $(x,\mu)$. $f_{1}(x,a)$ is once continuously differentiable and of at most quadratic growth in $(x,a)$.
(ii) $f_{0}(x,\mu)$ is convex in $x$ and there exists a positive constant $b_{4}>0$, such that
$|\partial_{x}f_{0}(x,\mu)-\partial_{x}f_{0}(x^{\prime},\mu^{\prime})|<b_{4}\left(|x-x^{\prime}|+W_{2}(\mu,\mu^{\prime})\right).$ (4.25)
(iii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds
$f_{1}(x^{\prime},a^{\prime})-f_{1}(x,a)-\partial_{(x,a)}f_{1}(x,a)\cdot(x^{\prime}-x,a^{\prime}-a)$ (4.26)
$\geq$ $\iota(x^{\prime}-x)^{2}+\eta(a^{\prime}-a)^{2}.$
(iv) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipchitz continuous in $(x,a)$.
(v) $A=\mathbb{R}$, and it holds that
$\min\left\{2\iota-\frac{\ell^{2}}{2\eta}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}-2b_{4},\frac{2b_{3}^{2}\eta}{\zeta^{2}}-\frac{b_{3}\ell}{2\eta}-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.27)
Remark 4.5
If we set $f_{0}(x,\mu)=b_{4}x\bar{\mu},f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C,\ell=0$. Then the requirement in (4.27) becomes
$\min\left\{2A-\frac{|b_{2}|}{2}-|b_{4}|,\frac{b_{3}^{2}}{2C}-\frac{|b_{2}|}{2}\right\}>\frac{r}{2}.$ (4.28)
Fixing $C,b_{2},b_{4},$ we take a large $b_{3}$ such that $\frac{b_{3}^{2}}{2C}-\frac{|b_{2}|}{2}$ is greater than $r/2$. Then we choose a sufficiently large $A$ such that $2A-\frac{|b_{2}|}{2}-|b_{4}|$ exceeds $r/2$. This construction satisfies all the required conditions.
Proof. It’s clear that Assumption 3.1 is satisfied. And by Lemma 4.3, $\hat{\alpha}$ is Lipchitz. In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+f_{0}(x,\mu)+f_{1}(x,a)$ is convex in $(x,a)$ according to assumption on $f_{0},f_{1}$, thus Assumption 4.1 is satisfied. To prove that the BSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$, the only condition that remains to be verified is (2.5).
We consider the FBSDE (4.12), and set
$B(t,x,y,\mu)=b_{1}x+b_{2}\overline{\mu}+b_{3}\hat{\alpha}(x,y),$ (4.29)
$F(t,x,y,\mu)=b_{1}y+\partial_{x}f_{0}(x,\mu)+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
Take four arbitrary square integrable random variables $X,Y,X^{\prime},Y^{\prime}$. Define $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}})$ . We have
$-r\hat{X}\hat{Y}-\hat{X}[F(t,U)-F(t,U^{\prime})]+\hat{Y}[B(t,U)-B(t,U^{\prime})]$ (4.30)
$=$ $-r\hat{X}\hat{Y}-\hat{X}\left((b_{1}-r)\hat{Y}+\partial_{x}f_{0}(X,\mathcal{L}_{X})-\partial_{x}f_{0}(X^{\prime},\mathcal{L}_{X^{\prime}})+\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
$+\hat{Y}\left(b_{1}\hat{X}+{b}_{2}\mathbb{E}[\hat{X}]+b_{3}(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
$\leq$ $-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)+b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$
$+b_{2}\hat{Y}\mathbb{E}[\hat{X}]+b_{4}|\hat{X}|\cdot\left(|\hat{X}|+(\mathbb{E}[\hat{X}^{2}])^{1/2}\right).$
Since $f_{1}$ is $\iota$-convex in $x$, we have
$\left[\partial_{x}f_{1}(x^{\prime},a)-\partial_{x}f_{1}(x,a)\right](x^{\prime}-x)\geq 2\iota(x^{\prime}-x)^{2}.$ (4.31)
Moreover, $\partial_{x}f_{1}$ is $l$-Lipschitz in $a$ and $\hat{\alpha}$ satisfies (4.22), we have that
$-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))\right)$ (4.32)
$=$ $-\hat{X}\left(\partial_{x}f_{1}(X,\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))\right)$
$-\hat{X}\left(\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X,Y))-\partial_{x}f_{1}(X^{\prime},\hat{\alpha}(X^{\prime},Y^{\prime}))\right)$
[…]
From Lemma 4.3, $\hat{\alpha}$ satisfies (4.24), it follows that
$b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$ (4.33)
$=$ $b_{3}\hat{Y}\left(\hat{\alpha}(X,Y)-\hat{\alpha}(X,Y^{\prime})\right)+b_{3}\hat{Y}\left(\hat{\alpha}(X,Y^{\prime})-\hat{\alpha}(X^{\prime},Y^{\prime})\right)$
[…]
By applying elementary estimates, we derive
$\mathbb{E}\left[-r\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (4.34)
$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2}+2b_{4})\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta}+\frac{|b_{2}|}{2})\mathbb{E}[\hat{Y}^{2}]$
$<$ $-\frac{r}{2}\mathbb{E}[\hat{X}^{2}+\hat{Y}^{2}].$
Now we know (4.12) admits a unique solution $X_{t}^{\xi}\in L^{2}_{r}$ and set $\bar{\mu}_{t}=\mathbb{E}[X_{t}^{\xi}]$ For the FBSDE(4.13), we set
$B^{\prime}(t,x,y)=b_{1}x+b_{2}\overline{\mu}_{t}+b_{3}\hat{\alpha}(x,y),$ (4.35)
$F^{\prime}(t,x,y)=b_{1}y+\partial_{x}f_{0}(x,\mu_{t})+\partial_{x}f_{1}(x,\hat{\alpha}(x,y))-ry.$
Following the identical analytical procedure, we have
$\mathbb{E}\left[-r\hat{X}\hat{Y}-\hat{X}(F^{\prime}(t,X,Y)-F^{\prime}(t,X^{\prime},Y^{\prime}))+\hat{Y}(B^{\prime}(t,X,Y)-B^{\prime}(t,X^{\prime},Y^{\prime}))\right]$ (4.36)
$\leq$ $(-2\iota+\frac{l^{2}}{2\eta}+\frac{b_{3}l}{2\eta}+b_{4})\mathbb{E}[\hat{X}^{2}]+(-\frac{2b_{3}^{2}\eta}{\zeta^{2}}+\frac{b_{3}l}{2\eta})\mathbb{E}[\hat{Y}^{2}]$
$<$ $-\frac{r}{2}\mathbb{E}[\hat{X}^{2}+\hat{Y}^{2}].$
5 Master equation representation
While we have derived a Nash equilibrium solution through FBSDEs that yields identical feedback forms for both the representative player and social equilibrium, this feedback structure differs from our previously defined formulation in Definition 3.2. In this section, we shall establish an alternative representation of the Nash equilibrium using classical solutions to the elliptic master equation (1.7).
Following Assumption 4.1, we define
$H(x,\mu,y)=H_{0}(x,\mu,\hat{\alpha}(x,y),y).$ (5.1)
Through the assumptions on $b,f$, we can easily deduce the relationship
$\partial_{y}H(x,\mu,y)=b(x,\mu,\hat{\alpha}(x,y)).$ (5.2)
Assume the master equation (1.7) has a classical solution $U(x,\mu)\in{\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ with $H$ defined above, and $\hat{b}(x,\mu)=b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$. For any $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$, suppose the following SDE admits a unique solution in $L^{2}_{r}$:
$\mathcal{X}_{t}^{\xi}=\xi+\int_{0}^{t}b\left(\mathcal{X}_{s}^{\xi},\mathcal{L}_{\mathcal{X}_{s}^{\xi}},\hat{\alpha}\left(\mathcal{X}_{s}^{\xi},\partial_{x}U\left(\mathcal{X}_{s}^{\xi},\mathcal{L}_{\mathcal{X}_{s}^{\xi}}\right)\right)\right)\mathrm{d}s+B_{t}.$ (5.3)
Then we take $\mathcal{Y}_{t}^{\xi}=\partial_{x}U(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}})$. By differentiating both sides of the master equation (1.7) with respect to $x$ and applying Itô’s formula, we obtain
$\begin{split}\mathrm{d}\mathcal{Y}_{t}^{\xi}=&\Bigg{[}\partial_{xx}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\cdot b\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\hat{\alpha}\left(\mathcal{X}_{t}^{\xi},\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\right)\right)\\ &+\frac{1}{2}\partial_{xx}\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)+\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\Bigg{[}\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\tilde{\mathcal{X}}^{\xi}_{t}\right)\\ &+\partial_{y}H\left(\tilde{\mathcal{X}}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\partial_{x}U\left(\tilde{\mathcal{X}}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\right)\cdot\partial_{\mu}\partial_{x}U\left(\mathcal{X}^{\xi}_{t},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\tilde{\mathcal{X}}^{\xi}_{t}\right)\Bigg{]}\Bigg{]}\mathrm{d}t\\ &+\partial_{xx}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\mathrm{d}B_{t}\\ =&\left(r\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)-\partial_{x}H\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}},\partial_{x}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\right)\right)\mathrm{d}t+\partial_{xx}U\left(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}}\right)\mathrm{d}B_{t}.\end{split}$ (5.4)
By comparing it with (4.12), we derive the following relationship for social equilibrium:
$Y_{t}^{\xi}=\partial_{x}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{\xi}=\partial_{xx}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}).$ (5.5)
Applying the same argument to (4.13), we obtain:
$Y_{t}^{x}=\partial_{x}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x}=\partial_{xx}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}).$ (5.6)
This demonstrates that both the representative player and social equilibrium employ the same closed-loop control
$\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu)).$ (5.7)
We now revisit the mean field games through the master equation. Let $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ be the initial state with distribution $\mu$. We prove that under the assumption that the master equation admits a classical solution, the feedback control defined by (5.7) constitutes a Nash equilibrium. Moreover, the solution to the master equation is precisely the value function of the representative player.
Theorem 5.1
Assume the master equation (1.7) admits a classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ which is of at most quadratic growth, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
$\begin{cases}X_{t}^{\xi}=\xi+\int_{0}^{t}b(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}},\hat{\alpha}(X_{s}^{\xi},\partial_{x}U(X_{s}^{\xi},\mathcal{L}_{X_{s}^{\xi}})))\mathrm{d}t+B_{t},\\ X_{t}^{x}=x+\int_{0}^{t}b(X_{s}^{x},\mathcal{L}_{X_{s}^{x}},\hat{\alpha}(X_{s}^{x},\partial_{x}U(X_{s}^{x},\mathcal{L}_{X_{s}^{x}})))\mathrm{d}t+B_{t}.\end{cases}$ (5.8)
Then for every initial distribution $\mu_{0}$, $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium, and $U(x,\mu_{0})$ is exactly the value function given $\alpha^{*}$.
Proof. Given $\alpha^{*}$ and initial $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu_{0})$, the states of social equilibrium is governed by the following SDE:
$X_{t}^{\xi}=\xi+\int_{0}^{t}\partial_{y}H(X_{s}^{\xi},\rho_{s},\partial_{x}U(X_{s}^{\xi},\rho_{s}))\mathrm{d}s+B_{t},\quad\quad\rho_{s}\triangleq\mathcal{L}_{X_{s}^{\xi}}.$ (5.9)
We note that this SDE is the same as (5.3). Meanwhile, the state of the representative player is governed by
$X_{t}^{x}=x+\int_{0}^{t}b(X_{s}^{x},\rho_{s},\beta_{s})\mathrm{d}s+B_{t},$ (5.10)
where $\beta\in\mathcal{A}$ remains to be determined. Applying Itô’s formula to $\mathrm{e}^{-rt}U(X^{x}_{t},\rho_{t})$ yields
$\begin{split}\mathbb{E}\mathrm{e}^{-rT}U(X^{x}_{T},\rho_{T})=&U(x,\mu_{0})+\mathbb{E}\int_{0}^{T}\big{[}-r\mathrm{e}^{-rt}U(X^{x}_{t},\rho_{t})+\mathrm{e}^{-rt}\partial_{x}U(X^{x}_{t},\rho_{t})\cdot b(X_{t}^{x},\rho_{t},\beta_{t})\\ &+\frac{1}{2}\mathrm{e}^{-rt}\partial_{xx}U(X^{x}_{t},\rho_{t})+\mathrm{e}^{-rt}\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\big{[}\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(X^{x}_{t},\rho_{t},\tilde{X}^{\xi}_{t})\\ &+\partial_{y}H(\tilde{X}_{t}^{\xi},\rho_{t},\partial_{x}U(\tilde{X}_{t}^{\xi},\rho_{t}))\cdot\partial_{\mu}U(X^{x}_{t},\rho_{t},\tilde{X}^{\xi}_{t})\big{]}\big{]}\mathrm{d}t\\ =&U(x,\mu_{0})+\mathbb{E}\int_{0}^{T}\mathrm{e}^{-rt}\left[\partial_{x}U(X^{x}_{t},\rho_{t})\cdot b(X_{t}^{x},\rho_{t},\beta_{t})-H(X^{x}_{t},\rho_{t},\partial_{x}U(X^{x}_{t},\rho_{t}))\right]\mathrm{d}t\\ \geq&U(x,\mu_{0})-\mathbb{E}\int_{0}^{T}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\beta_{t})\mathrm{d}t.\end{split}$ (5.11)
Since $U$ is of at most quadratic growth, taking the limit as $T\rightarrow+\infty$ yields
$U(x,\mu_{0})\leq\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\beta_{t})\mathrm{d}t$ (5.12)
for every feasible control $\beta$, and the equality holds when $\beta_{t}=\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t}))$. This shows $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is an Nash equilibrium. And we have
$U(x,\mu_{0})=\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t})))\mathrm{d}t,$ (5.13)
this completes the proofs of our desired results.
At the end of this section, we provide an example of a solvable elliptic master equation. Set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a,f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, where $b_{1},b_{2},b_{3},b_{4},A,C$ are constants satisfying all conditions in Theorem 4.4 and Remark 4.5. Then we have $\hat{\alpha}(x,y)=-\frac{b_{3}y}{2C}$ and
$H(x,\mu,y)=(b_{1}x+b_{2}\bar{\mu})\cdot y+b_{4}x\bar{\mu}+Ax^{2}-\frac{b_{3}^{2}}{4C}y^{2}.$ (5.14)
The master equation (1.7) now becomes
$\begin{split}rU(x,\mu)=&(b_{1}x+b_{2}\bar{\mu})\cdot\partial_{x}U(x,\mu)+b_{4}x\bar{\mu}+Ax^{2}\\ &-\frac{b_{3}^{2}}{4C}(\partial_{x}U(x,\mu))^{2}+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})(b_{1}\tilde{\xi}+b_{2}\bar{\mu}-\frac{b^{2}_{3}}{2C}\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (5.15)
Theorem 5.2
The master equation (5.15) admits a classical solution $U(x,\mu)$ in ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$.
Proof. We assume that the solution takes the form
$U(x,\mu)=a_{1}x^{2}+a_{2}x\bar{\mu}+a_{3}(\bar{\mu})^{2}+a_{4},$ (5.16)
where $a_{1},a_{2},a_{3},a_{4}$ are constants remain to be determined. Then we have
$\begin{split}&\partial_{x}U(x,\mu)=2a_{1}x+a_{2}\bar{\mu},\quad\quad\quad\partial_{xx}U(x,\mu)=2a_{1},\\ &\partial_{\mu}U(x,\mu,\tilde{x})=a_{2}x+2a_{3}\bar{\mu},\quad\quad\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{x})=0.\end{split}$ (5.17)
Substituting these into the equation (5.15), we obtain
$\begin{split}&r\left(a_{1}x^{2}+a_{2}x\bar{\mu}+a_{3}(\bar{\mu})^{2}+a_{4}\right)\\ =&(b_{1}x+b_{2}\bar{\mu})\cdot(2a_{1}x+a_{2}\bar{\mu})+b_{4}x\bar{\mu}+Ax^{2}-\frac{b_{3}^{2}}{4C}(2a_{1}x+a_{2}\bar{\mu})^{2}+a_{1}\\ &+(a_{2}x+2a_{3}\bar{\mu})\cdot\left(b_{1}\bar{\mu}+b_{2}\bar{\mu}-\frac{b^{2}_{3}}{2C}(2a_{1}\bar{\mu}+a_{2}\bar{\mu})\right).\end{split}$ (5.18)
Comparing all coefficients, we get the following system of linear and quadratic equations:
$\begin{cases}ra_{1}=2b_{1}a_{1}+A-\frac{b_{3}^{2}}{C}a_{1}^{2},\\ ra_{2}=2b_{2}a_{1}+b_{1}a_{2}+b_{4}-\frac{b_{3}^{2}}{C}a_{1}a_{2}+a_{2}(b_{1}+b_{2}-\frac{b_{3}^{2}}{C}a_{1}-\frac{b_{3}^{2}}{2C}a_{2}),\\ ra_{3}=b_{2}a_{2}-\frac{b_{3}^{2}}{4C}a_{2}^{2}+2a_{3}(b_{1}+b_{2}-\frac{b_{3}^{2}}{C}a_{1}-\frac{b_{3}^{2}}{2C}a_{2}),\\ ra_{4}=a_{1}.\end{cases}$ (5.19)
Then, we can solve for $a_{1},a_{2},a_{3},a_{4}$.
Remark 5.3
We have
$\hat{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))=(b_{1}-\frac{a_{1}b_{3}^{2}}{C})x+(b_{2}-\frac{a_{2}b_{3}^{2}}{2C})\bar{\mu}.$ (5.20)
Since $b_{1}+|b_{2}|<\frac{r}{2}$, if we further require $a_{1}>0,a_{2}>0$, the SDE (5.3) admits a unique solution in $L_{r}^{2}$.
6 Viscosity solution to distribution-dependent elliptic PDE
We have proven that the classical solutions of the master equations can be employed to resolve the infinite-time FBSDEs. In this section, the solution of the infinite-time FBSDEs will be utilized to characterize the viscosity solutions of the distribution-dependent elliptic PDEs. Specifically, we consider the following FBSDEs with initial state $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $x\in\mathbb{R}$:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\partial_{y}H(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[\partial_{x}{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})-rY_{t}^{\xi}\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.1)
$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (6.2)
Here $H(x,\mu,y)$ is defined in (5.1) and the above equations are the same as (1.5) and (1.6).
We define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ and attempt to prove that it satisfy the following elliptic PDE:
$\begin{split}r\mathcal{U}(x,\mu)=&\partial_{x}H(x,\mu,\mathcal{U}(x,\mu))+\partial_{y}H(x,\mu,\mathcal{U}(x,\mu))\cdot\partial_{x}\mathcal{U}(x,\mu)+\frac{1}{2}\partial_{xx}\mathcal{U}(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})+\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\mathcal{U}(\tilde{\xi},\mu))\right].\end{split}$ (6.3)
This equation is derived by taking the partial derivative of both sides of the master equation (1.7) with respect to $x$. We have the relationship $\mathcal{U}(x,\mu)=\partial_{x}U(x,\mu)$.
The proof of $\mathcal{V}$ being a viscosity solution to equation (6.3) rests on two fundamental results:
  • The value of $Y^{x,\xi}_{0}$ depends solely on the distribution of $\xi$, but not on the specific realization of $\xi$.
  • $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$.
6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
We shall prove the weak uniqueness for the infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.4)
We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}$. We then give the definitions of strong and weak uniqueness.
Definition 6.1 (Strong uniqueness)
We say that the strong uniqueness holds for FBSDE (6.4) if on any admissible set-up $(\Omega,{\cal F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B)$, for any two $\mathbb{F}$-progressively measurable $L^{2}_{r}$ three-tuples
$(X^{1}_{t},Y^{1}_{t},Z^{1}_{t})_{t\geq 0},\quad(X^{2}_{t},Y^{2}_{t},Z^{2}_{t})_{t\geq 0}$ (6.5)
satisfying the FBSDE (6.4) with the same initial condition $X_{0}$ (up to an exceptional event), it holds that
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|X^{1}_{t}-X^{2}_{t}|^{2}+|Y^{1}_{t}-Y^{2}_{t}|^{2}+|Z^{1}_{t}-Z^{2}_{t}|^{2}\right)\mathrm{d}t\right]=0$ (6.6)
Definition 6.2 (Weak uniqueness)
For any two set-ups $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1},\mathbb{F}^{1})$ and $(\Omega^{2},\mathcal{F}^{2},\mathbb{P}^{2},\mathbb{F}^{2})$ with inputs $(X_{0}^{1},B^{1})$ and $(X_{0}^{2},B^{2})$, $X_{0}^{1},X_{0}^{2}$ have the same law on $\mathbb{R}$. We say the weak uniqueness holds for FBSDE (6.4) if the $L^{2}_{r}$ solutions $(X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on corresponding set-ups have the same distribution.
We use the same scheme as the one developed by Yamada and Watanabe to prove that path-wise uniqueness of solutions of FBSDE’s implies uniqueness in the sense of probability law.
Theorem 6.3
Assume that on $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B_{t})_{t\geq 0}$, the FBSDE has a unique strong solution $(X_{t},Y_{t},Z_{t})$. Then the law of $(X_{0},B_{t},X_{t},Y_{t},\int_{0}^{t}Z_{s}\mathrm{d}s)$ only depends on $\mathcal{L}(X_{0})$.
Proof. At first we denote by $C([0,\infty);\mathbb{R})$ the space of continuous $\mathbb{R}$-valued functions on $0,\infty)$ equipped with the metric of uniform convergence on compacts:
$d(\omega^{1},\omega^{2})=\sum_{n\geq 0}2^{-n}\sup_{t\in[0,n]}\max(|\omega_{t}^{1}-\omega_{t}^{2}|,1).$ (6.7)
And define
$\begin{split}&\Omega_{\text{input}}\triangleq\mathbb{R}\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{output}}\triangleq C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{canon}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}.\end{split}$ (6.8)
Let us consider two filtered probability spaces $(\Omega^{i},\mathcal{F}^{i},\mathbb{P}^{i},\mathbb{F}^{i}),i=1,2$. Denote by $Q^{1}$ and $Q^{2}$ the distribution of $(X_{0}^{1},B_{t}^{1},X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{0}^{2},B_{t}^{2},X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on $\Omega_{\text{canon}}=\Omega_{\text{input}}\times\Omega_{\text{output}}$, by $Q_{\text{input}}$ the common distribution of the processes $(X_{0}^{1},B_{t}^{1}),(X_{0}^{2},B_{t}^{2})$ on $\Omega_{\text{input}}$.
Let us now define for $i\in\{1,2\}$,
$q^{i}(\omega_{\text{input}};F):\Omega_{\text{input}}\times\mathcal{B}(\Omega_{\text{output}})\rightarrow[0,1]$ (6.9)
as the regular conditional probability for $\mathcal{B}(\Omega_{\text{output}})$ given $\omega_{\text{input}}\in\Omega_{\text{input}}$ (under $Q^{i}$). It satisfies:
  • $\forall\omega_{\text{input}}\in\Omega_{\text{input}}$, $q^{i}(\omega_{\text{input}};\cdot)$ is a probability measure on $(\Omega_{\text{output}},\mathcal{B}(\Omega_{\text{output}}))$.
  • $\forall F\in\mathcal{B}(\Omega_{\text{output}})$, the mapping $(\xi,w)\rightarrow q^{i}(\xi,w;F)$ is ${\cal B}(\mathbb{R})\otimes{\cal B}(C([0,\infty),\mathbb{R}))$-measurable.
  • $\forall F\in{\cal B}(\Omega_{\text{output}}),\forall G\in{\cal B}(\Omega_{\text{input}})$:
    $Q^{i}(G\times F)=\int_{G}q^{i}(\omega_{\text{input}};F)Q_{\text{input}}(\mathrm{d}\omega_{\text{input}}).$ (6.10)
Next, we need a enlarged space $(\Omega_{\text{total}},\mathcal{G},Q)$ to support all processes. Define
$\Omega_{\text{total}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}\times\Omega_{\text{output}},$ (6.11)
and $\mathcal{G}$ is the completion of the $\sigma$-field ${\cal B}(\Omega_{\text{canon}})\otimes{\cal B}(\Omega_{\text{output}})$ by the collection $\cal N$} of all null sets under the probability measure
$Q(G\times F_{1}\times F_{2})=\int_{G}q^{1}(\omega_{\text{input}};F_{1})q^{2}(\omega_{\text{input}};F_{2})Q_{\text{input}}(\mathrm{d}\omega_{\text{input}}),$ (6.12)
where $F_{1},F_{2}\in{\cal B}(\Omega_{\text{output}}),G\in{\cal B}(\Omega_{\text{input}})$.
We observe that $Q(G\times F_{1}\times\Omega_{\text{output}})=Q^{1}(G\times F_{1}),Q(G\times\Omega_{\text{output}}\times F_{2})=Q^{2}(G\times F_{2}).$ Inparticular, we denote by $(\xi,w,x^{1},y^{1},\zeta^{1},x^{2},y^{2},\zeta^{2})$ the canonical process on $\Omega_{\text{total}}$, then $(\xi,w,x^{1},y^{1},\zeta^{1})$ has distribution $Q^{1}$ and $(\xi,w,x^{2},y^{2},\zeta^{2})$ has distribution $Q^{2}$.
We define $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ by
$z^{i}_{t}=\begin{cases}\lim_{n\rightarrow\infty}n\left(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+}\right)\quad~\mbox{if the limit exists,}\\ 0\quad~\mbox{otherwise.}\end{cases}$ (6.13)
[…]
$\zeta^{i}_{t}=\int_{0}^{t}z^{i}_{s}\mathrm{d}s,\quad t\geq 0.$ (6.14)
Moreover, we have
$\begin{split}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|z^{i}_{t}|\mathrm{d}t&\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|n(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+})|^{2}\mathrm{d}t\\ &=\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n^{2}\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}|\int_{(t-\frac{1}{n})+}^{t}Z_{s}^{i}\mathrm{d}s|^{2}\mathrm{d}t\\ &\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}\int_{(t-\frac{1}{n})+}^{t}|Z_{s}^{i}|^{2}\mathrm{d}s\mathrm{d}t\\ &\leq\mathbb{E}^{Q^{i}}\int_{0}^{\infty}\mathrm{e}^{rs}|Z_{s}^{i}|^{2}\mathrm{d}s,\end{split}$ (6.15)
the last inequality following from Fubini’s theorem.
Let us now endow $(\Omega_{\text{total}},{\cal G},Q)$ with the filtration $\mathbb{G}$, where $\mathbb{G}=\{{\cal G}_{t}\}_{t\geq 0}$ is the complete and right-continuous augumentation under $Q$ of the canonical filtration
${\cal H}_{t}=\{(\xi,w_{s},x^{1}_{s},y^{1}_{s},\zeta^{1}_{s},x^{2}_{s},y^{2}_{s},\zeta^{2}_{s});0\leq s\leq t\}$ (6.16)
on $\Omega_{\text{total}}$. It’s easy to see that $\xi$ is ${\cal G}_{0}$-measurable and that $(w_{s},x^{1}_{s},y^{1}_{s},z^{1}_{s},x^{2}_{s},y^{2}_{s},z^{2}_{s})_{t\geq 0}$ is $\{{\cal G}_{t}\}_{t\geq 0}$-progressively measurable. Moreover, for $i\in\{1,2\}$,
$Q\left\{\omega\in\Omega_{\text{total}};(x,w,x^{i},y^{i},\zeta^{i})\in A\right\}=Q^{i}\left\{(X_{0}^{i},B^{i},X^{i},Y^{i},\int_{0}^{\cdot}Z^{i}_{s}\mathrm{d}s)\in A\right\};\quad A\in{\cal B}(\Omega_{\text{canon}}).$ (6.17)
Actually, we just have to prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion: we follow the proof given in ([6], Remark 1.6). Let us firstly define
$\pi_{t}:C([0,\infty);\mathbb{R})\rightarrow C([0,t];\mathbb{R}),\quad h\rightarrow h|_{[0,t]},$ (6.18)
and
$\pi^{\prime}_{t}:\gamma\rightarrow\gamma_{t},\quad h\rightarrow h|_{[0,t]},$ (6.19)
where $\gamma=\Omega_{\text{output}}$ and
$\gamma_{t}=C([0,t];\mathbb{R})\times C([0,t];\mathbb{R})\times C([0,t];\mathbb{R}).$ (6.20)
Endowing $C([0,t];\mathbb{R})$ and $\gamma_{t}$ with their borelian $\sigma$-fields, we define
${\cal K}_{t}\triangleq\sigma\{\pi_{t}\},\quad{\cal K}^{\prime}_{t}\triangleq\sigma\{\pi^{\prime}_{t}\}.$ (6.21)
Using the separability of the spaces $C([0,t];\mathbb{R})$ and $\gamma_{t}$, we see that
${\cal K}_{t}=\sigma\{w_{s};0\leq s\leq t\},$ (6.22)
and that $\forall i\in\{1,2\},\forall A\in{\cal K}^{\prime}_{t}$, the set $\{(X^{i},Y^{i},\int_{0}^{\cdot}Z^{i}_{s}\mathrm{d}s)\in A\}$ belongs to ${\cal F}^{i}_{t}$.
Now, considering $A\in{\cal K}^{\prime}_{t}$ , we want to show that, for $i\in\{1,2\}$, the map
$\Omega_{\text{input}}\rightarrow[0,1],\quad(\xi,w)\rightarrow q^{i}(\xi,w;A)$ (6.23)
is measurable with respect to the completion of the $\sigma$-field ${\cal B}(\mathbb{R})\otimes{\cal K}_{t}$ under the probability measure $Q_{\text{input}}$, denoted $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$. Indeed, let us consider $F\in{\cal B}(\mathbb{R})$, $G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$. Then, $\forall i\in\{1,2\}$
$\begin{split}&\int I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)q^{i}(\xi,w;A)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\mathbb{E}^{\mathbb{P}^{i}}\left[I_{F}(X_{0})I_{G_{1}}(B^{i})I_{G_{2}}(B^{i})I_{A}(X^{i},Y^{i},\int_{0}^{\cdot}Z_{s}^{i}\mathrm{d}s)\right]\\ =&\mathbb{E}^{\mathbb{P}^{i}}\left[I_{F}(X_{0})I_{G_{1}}(B^{i})I_{A}(X^{i},Y^{i},\int_{0}^{\cdot}Z_{s}^{i}\mathrm{d}s)\right]\mathbb{E}^{\mathbb{P}^{i}}\left[I_{G_{2}}(B^{i})\right]\\ =&\int I_{F}(\xi)I_{G_{1}}(w)q^{i}(\xi,w;A)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\int I_{G_{2}}(w)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w).\end{split}$ (6.24)
Hence, the map $(\xi,w)\rightarrow q^{i}(\xi,w;A)$ is measurable with respect to $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$.
Now we prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion. Let us consider $(A,A^{\prime})\in({\cal K}^{\prime}_{t})^{2},F\in{\cal B}(\mathbb{R}),G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$. Then,
$\begin{split}&\mathbb{E}^{Q}\left[I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)I_{A}(x^{1},y^{1},\zeta^{1})I_{A^{\prime}}(x^{2},y^{2},\zeta^{2})\right]\\ =&\int I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)q^{1}(\xi,w;A)q^{2}(\xi,w;A^{\prime})Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\int I_{F}(\xi)I_{G_{1}}(w)q^{1}(\xi,w;A)q^{2}(\xi,w;A^{\prime})Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\int I_{G_{2}}(w)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\mathbb{E}^{Q}\left[I_{F}(\xi)I_{G_{1}}(w)I_{A}(x^{1},y^{1},\zeta^{1})I_{A^{\prime}}(x^{2},y^{2},\zeta^{2})\right]\mathbb{E}^{Q}\left[I_{G_{2}}(w)\right].\end{split}$ (6.25)
Noting that ${\cal H}^{t}={\cal B}(\mathbb{R})\otimes{\cal K}_{t}\otimes{\cal K}^{\prime}_{t}\otimes{\cal K}^{\prime}_{t}$, we conclude that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion.
At last, applying the same procedure on $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in ([5], Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
$\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}B(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.26)
Through the strong uniqueness, we know that under $Q$, the processes $(x^{1}_{t},y^{1}_{t},\zeta^{1}_{t})_{t\geq 0}$ and $(x^{2}_{t},y^{2}_{t},\zeta^{2}_{t})_{t\geq 0}$ have the same law, then we get the desired result.
Remark 6.4
For any two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution and $x\in\mathbb{R}$, the solutions $Y^{x,\xi_{1}},Y^{x,\xi_{2}}$ to (6.2) are the same. We can say that $Y^{x,\xi}_{0}$ depends solely on the distribution of $\xi$.
6.2 Regularity of the FBSDE solutions
Lemma 6.5
For FBSDEs (6.1) and (6.2), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$. Assuming all assumptions in Theorem 4.4 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
$|\mathcal{V}(x,\mu)-\mathcal{V}(x^{\prime},\mu^{\prime})\leq C(|x-x^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime}))$ (6.27)
Proof. Let $(X^{\xi_{1}},Y^{\xi_{1}},X^{x_{1},\xi_{1}},Y^{x_{1},\xi_{1}})$ and $(X^{\xi_{2}},Y^{\xi_{2}},X^{x_{2},\xi_{2}},Y^{x_{2},\xi_{2}})$ be the solutions of Equations (6.1) and (6.2) with $\mathcal{L}_{\xi_{1}}=\mu_{1},\mathcal{L}_{\xi_{2}}=\mu_{2}$. Set
$\begin{split}&\hat{X}^{\xi}=X^{\xi_{1}}-X^{\xi_{2}},\quad\hat{Y}^{\xi}=Y^{\xi_{1}}-Y^{\xi_{2}},\\ &\hat{X}^{x,\xi}=X^{x_{1},\xi_{1}}-X^{x_{2}\xi_{2}},\quad\hat{Y}^{x,\xi}=Y^{x_{1},\xi_{1}}-Y^{x_{2},\xi_{2}}.\end{split}$ (6.28)
$C_{1},C_{2},C_{3},C_{4}$ appeared in the following proof are positive constants.
Applying Itô’s formula to $\mathrm{e}^{-rt}|Y^{x_{1},\xi_{1}}_{t}-Y^{x_{2},\xi_{2}}_{t}|^{2}$, we get
$\begin{split}|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}&\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+W_{2}^{2}(X^{\xi_{1}}_{t},X^{\xi_{2}}_{t})\right]\mathrm{d}t\\ &\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t.\end{split}$ (6.29)
Applying Itô’s formula to $\mathrm{e}^{-rt}\hat{X}^{x,\xi}\hat{Y}^{x,\xi}$, we get
$\begin{split}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}\right]\mathrm{d}t\leq C_{2}\left(\hat{X}^{x,\xi}_{0}\hat{Y}^{x,\xi}_{0}+\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\right).\end{split}$ (6.30)
It turns out that
$\begin{split}(\hat{Y}^{x,\xi}_{0})^{2}\leq&C_{1}C_{2}\hat{X}^{x,\xi}_{0}\hat{Y}^{x,\xi}_{0}+(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\\ \leq&\frac{1}{2}(\hat{Y}^{x,\xi}_{0})^{2}+\frac{C_{1}^{2}C_{2}^{2}}{2}(\hat{X}^{x,\xi}_{0})^{2}+(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t.\end{split}$ (6.31)
Thus
$(\hat{Y}^{x,\xi}_{0})^{2}\leq C_{1}^{2}C_{2}^{2}(\hat{X}^{x,\xi}_{0})^{2}+2(C_{1}C_{2}+C_{1})\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t.$ (6.32)
Applying the same arguments to $\mathrm{e}^{-rt}|\hat{Y}^{\xi}_{t}|^{2}$ and $\mathrm{e}^{-rt}\hat{X}^{\xi}\hat{Y}^{\xi}$, we can get that
$\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}(\hat{X}^{\xi}_{t})^{2}\mathrm{d}t\leq C_{3}\mathbb{E}|\xi_{1}-\xi_{2}|^{2}.$ (6.33)
So we can deduce that
$|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathbb{E}|\xi_{1}-\xi_{2}|^{2}\right).$ (6.34)
Since $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$, we have
$|\mathcal{V}(x_{1},\mu_{1})-\mathcal{V}(x_{2},\mu_{2})|^{2}\leq C_{4}\left(|x_{1}-x_{2}|^{2}+\mathcal{W}_{2}^{2}(\mu_{1},\mu_{2})\right),$ (6.35)
then we get the desired result.
6.3 Connection with distribution-dependent elliptic PDE
Now let us give the definition of a viscosity solution for PDE (6.3). We rewrite the PDE as follows:
$(\mathcal{L}\mathcal{U})(x,\mu,\mathcal{U}(x,\mu))+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.36)
where
$\begin{split}(\mathcal{L}\Phi)(x,\mu,\Psi)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.37)
and
$F(x,\mu,\Psi)\triangleq\partial_{x}H(x,\mu,\Psi(x,\mu))-r\Psi(x,\mu).$ (6.38)
Definition 6.6
Let $\mathcal{U}\in C(\mathbb{R}\times\mathcal{P}_{2})$. Then $\mathcal{U}$ is called a viscosity subsolution (resp. supersolution) of PDE (6.36) if, whenever $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp. minimum) of $\mathcal{U}-\Psi$, we have
$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.39)
(respectively
$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.40)
). The function $\mathcal{U}$ is called a viscosity solution of PDE (6.36) if it is both a viscosity subsolution and a viscosity supersolution.
We now assert that
Theorem 6.7
Assume that both FBSDEs (6.1) and (6.2) admit unique solutions in $L^{2}_{r}$, $Y_{0}^{x,\xi}$ depends solely on the distribution of $\xi$. Then we define the function $\mathcal{V}(x,\mu)\triangleq Y_{0}^{x,\xi}$ with $\mathcal{L}_{\xi}=\mu$, and assume $\mathcal{V}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$. The function $\mathcal{V}(x,\mu)$ is a viscosity solution of PDE (6.36).
Proof. We only show that $\mathcal{V}$ is a viscosity subsolution of PDE (6.36). A similar argument will show that it is also a viscosity supersolution of (6.36).
Due to the uniqueness for solutions of FBSDE, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$. Let $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum of $\mathcal{U}-\Psi$. We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$. We suppose that
$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))<0,$ (6.41)
It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$, such that for all $(x,\mu)\in O$,
$\begin{cases}\mathcal{V}(x,\mu)\leq\Psi(x,\mu),\\ (\mathcal{L}\Psi)(x,\mu,\mathcal{V}(x,\mu))+F(x,\mu,\mathcal{V}(x,\mu))<0.\end{cases}$ (6.42)
Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2). For some $T>0$, let $\tau$ denote the stopping time
$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\mathcal{L}_{X_{t}^{\xi^{0}}})\in O\}\land T.$ (6.43)
We first note that the pair of processes
$(\bar{Y}\left(t\right),\bar{Z}\left(t\right))\triangleq(Y^{x^{0},\xi^{0}}_{t\wedge\tau},I_{\left[0,\tau\right]}(t)Z^{x^{0},\xi^{0}}_{t}),\quad 0\leq t\leq T,$ (6.44)
is the solution of BSDE
$\begin{split}\bar{Y}_{t}=&\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.45)
Next, it follows from Itô’s formula that the pair of processes
$(\hat{Y}_{t},\hat{Z}_{t})\triangleq(\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}),I_{\left[0,\tau\right]}(t)\partial_{x}\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}))$ (6.46)
is the solution of BSDE
$\begin{split}\hat{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\int_{t}^{T}I_{[0,\tau]}(s)\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\hat{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.47)
Define
$\beta_{s}=-\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))-F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))$ (6.48)
and
$\left(\tilde{Y}\left(t\right),\tilde{Z}\left(t\right)\right)=\left(\hat{Y}\left(t\right)-\bar{Y}\left(t\right),\hat{Z}\left(t\right)-\bar{Z}\left(t\right)\right).$ (6.49)
We have
$\begin{split}\tilde{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)\beta_{s}\mathrm{d}s\\ &-\int_{t}^{T}\tilde{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.50)
Therefore
$\tilde{Y}_{0}=\mathbb{E}\left[\tilde{Y}_{\tau}+\int_{0}^{\tau}\beta_{s}\mathrm{d}s\right]$ (6.51)
Now from the choice of $O$ and $\tau$, a.s.
$\tilde{Y}_{\tau}=\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})\geq 0,\quad\beta_{s}>0,\quad s\in[0,\tau].$ (6.52)
Consequently, $\tilde{Y}_{0}=\Psi(x^{0},\mu^{0})-\mathcal{V}(x^{0},\mu^{0})>0$, which contradicts the earlier assumption.